OpenStudy (anonymous):

for f(x)=1/x-5 and g(x)=x^2+2, find: (fog)(x) and (gof)(6)

OpenStudy (zzr0ck3r):

\[(f\circ g)(x)=f(g(x))=\frac{1}{g(x)-5}=\frac{1}{(x^2+2)-5}=\frac{1}{x^2-3}\] Do something similar for \((g\circ f(x)\)).

OpenStudy (anonymous):

how did you do it?

OpenStudy (shamil98):

(f o g)(x) = you plug in what g(x) is = to into the equation for f(x)

OpenStudy (shamil98):

if f(x) = 2x and g(x) = -2 (f o g)(x) = 2(-2) understand?..

OpenStudy (shamil98):

(g o f)(6) simply means you evaluate f(6) and then put that result into the equation for g(x).

OpenStudy (anonymous):

Okay i'm getting it. It would be 1/x^2-3, but what about the second one i'm even more confused.

OpenStudy (shamil98):

f(x)=1/x-5 First, find f(6)

OpenStudy (shamil98):

To start off the second problem, can you find f(6)?

OpenStudy (anonymous):

I have no idea :/

OpenStudy (shamil98):

f(6) means you substitute for all values of x. f(x) = 1/x-5 f(6) = 1/6-5 = 1/1 = 1 Understand?

OpenStudy (anonymous):

Okay

OpenStudy (shamil98):

From there, you plug in f(6) into g(x) g(x)=x^2+2 (g o f)(6) = 1^2 + 2 = 3

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

so 2 is (g o f)(6) = 1^2 + 2 and 1 is f(x)=1/x-5

OpenStudy (shamil98):

number 1 is (g o f) as the zzrockr said in his explanation it is 1/x^2-3

OpenStudy (ranga):

(g o f)(6) = g(f(6)) Find f(6) first. f(x) = 1/(x - 5) f(6) = 1/(6 - 5) = 1/1 = 1 g(f(6)) = g(1) g(x) = x^2 + 2 g(1) = 1^2 + 2 = 1 + 2 = 3

OpenStudy (ranga):

For the first part: If (f o g)(6) = f(g(6)) Find g(6) first: g(x) = x^2 + 2 g(6) = 6^2 + 2 = 36 + 2 = 38 f(g(6)) = f(38) f(x) = 1/(x - 5) f(38) = 1/ (38 - 5) = 1/33